# The gauge symmetry violation in the QCD ground state

### v3

@@ 6 Particular solution to QCD
From above and now on in this section we denote the corresponding full gluon propagator $D$ by $D^{PT}$ (for the explanation see the text below). Obviously, the expression for the PT full gluon propagator (6.38) describes the propagation of the PT massless gluons, since it has the PT singularity on the mass-shell $q^{2}=0$, i.e., the singularity of the free gluon propagator $\sim 1/q^{2}$. Let us note in advance that the equality $\xi=\xi_{0}=\lambda^{-1}$ takes place only for the regularized massless gluon fields. For their renormalized counterparts these gauge-fixing parameters differ by the corresponding renormalization constant.
The system of the regularized eqs.(6.38)-(6.41), which is uniquely defined, is free of all the types of the scale parameters having the dimensions of a mass squared, forbidden by the exact gauge symmetry of the QCD Lagrangian. Therefore, it constitutes that the gauge symmetry of the QCD ground state, reflected by this system, coincides with the symmetry of its Lagrangian. Such coincidence is analogous to QED where the abelian $U(1)$ gauge symmetry of the Lagrangian is the same as of its ground state (omitting formally the YM degrees of freedom here, this coincidence has been proven just in this way in [10]). However, it is necessary to remind that the coupling constant remains strong, so to call it as $D^{PT}$ is only a mere convention, that is why the particular solution to QCD can be called as conventional QCD as well. Just this conventional solution to QCD has been investigated in full details the last five decades since its formulation in [1].
The essential feature of the above-mentioned coincidence is the equality $\xi=\xi_{0}$. In this connection, let us note that one can start from the expression (3.8). It is the general decomposition of the full gluon propagator, depending on the one variable, into the independent tensor structures. Then it will automatically satisfy to the relation (3.7), called the ST identity, determining thus the gauge-fixing parameter. This has nothing to do with the statement that the ST identity is a consequence of the exact gauge symmetry. It becomes its consequence, indeed, when $\xi$ is fixed by the relation $\xi=\xi_{0}=\lambda^{-1}$. The well-known PT linear relation $\xi=\xi_{0}$ is the consequence of the exact gauge invariance both for the QCD Lagrangian and its ground state. In principle, the PT QCD or, equivalently, the conventional QCD is a well-elaborated quantum gauge field theory [2, 3, 4, 5, 6, 7, 8, 9], and for its framework of the corresponding dynamical equations of motion see recent review [34] (and references therein). However, this theory is not our concern (but for some clarifying discussion see appendix B). We present it here within our formalism for the readers convenience to directly compare it with the general solution to QCD.
For this aim it is useful to reproduce the system of the relations (5.36)-(5.37) for the PT QCD. It will look like
@@
The first equalities in these relations coincide with the corresponding equalities in the relations (5.36)-(5.37), i.e., in both theories the general and the PT QCD or, equivalently, the particular QCD, they are exact mathematical results. At the same time, the second equalities in these relations remain prescriptions, since they have to be put zero by hand, as emphasized above.
The dimensional regularization method (DRM) [4, 6, 27, 28, 29], formulated within the framework of the conventional QCD, provides a gauge-invariant scheme to correctly calculate the finite parts of the generally QD loop integrals, while omitting their QD constants, i.e., simply to ignore them and to deal further only with the logarithmic divergences of the PT. This was a prescription rather than an exact result, as pointed out in [4]. However, from now on this prescription has been put on a firm mathematical ground. How to understand such kind of equalities has been explained in section 5.
7 Discussion
Let us now discuss some interesting physical consequences of the general solution to QCD. It requires the explicit presence of the tadpole term in the QCD ground state at the fundamental quark-gluon level, i.e., it cannot be discarded in the QCD/YM theory by any means. This was exactly proven in section 4 and used in section 5. Then it will drastically change the structure of the full gluon propagator in the deep IR region $q^{2}\rightarrow 0$. For the sake of transparency and clarity let us use the free gluon propagator in the t’ Hooft-Feynman gauge $\lambda^{-1}=1$, i.e., $D^{0}_{\mu\nu}(q)=i{\delta_{\mu\nu}/q^{2}}$. Then the full gluon propagator (5.28), on account of the definitions (5.29) and (5.30), up to the first iteration term looks like
@@ 7 Discussion
where the terms of order $O_{\mu\nu}(1/q^{2})$ are not shown, for simplicity. In this iteration expansion is explicitly shown only singular part of the corresponding Laurent series, and thus it represents the summation of all the severe IR singularities (more singular than the PT IR singularity $\sim 1/q^{2}$), which can be accumulated by the full gluon propagator. Evidently, such a summation is only possible due to the existence of the tadpole term. The dependence of the coefficients $\Phi_{k}(g^{2},...)$ on the coupling constant is not known, i.e., such kind of series are the cluster expansions, mentioned above [23]. These coefficients, being the sum of the infinite number of terms by themselves, are resulted from the skeleton loop integrals, contributing into full gluon self-energy through the invariant function $\Pi(q^{2};D)$ because of the NL character of the iteration procedure. Evidently, the iteration ab  infinity leads to the essential singularity at $q^{2}\rightarrow 0$! in the full gluon propagator (7.44), which will remain massless, but depending on the tadpole term. Let us note that the same series can be obtained by expanding the initial NL transcendental relation (5.30) in powers of the ratio $(\Delta^{2}_{t}(D)/q^{2})$.
@@
From the NL iteration expressions (7.44)-(7.45) it follows that the full gluon propagator in the IR region ($q^{2}\rightarrow 0$) is dominated by the tadpole contribution, while in the UV region ($q^{2}\rightarrow\infty$) it will be dominated by the free gluon one. This is completely different to QED, where the dressed photon propagator behaves at small photon momentum as its free counterpart ($\sim 1/q^{2}$). This is also true for the full gluon propagator in the PT QCD with unbroken gauge symmetry, investigated in section 6. The principal difference between these theories is due to the existence of the tadpole term in the QCD ground state, which precisely controls the structure of a single full gluon propagator at small $q^{2}$. That is why the dressed PT gluon (6.38) and free gluon states can not appear in the physical spectrum at large distances ($q^{2}\rightarrow 0$), i.e., they will be suppressed in this region by the tadpole term. If this statement will survive the corresponding renormalization program beyond the PT then confinement of the PT and free gluon states will be explained. This is important since there is no any other dynamical mechanism to confine these gluon states to the vacuum of QCD or to keep them inside hadrons. To prove the confinement of the singular full gluon propagator itself due to the above-mentioned essential singularity, though possible, but requires completely different investigation within the theory of distributions [26], and thus is the subject of the future work. Let us note that by the NP program we mean that the full gluon propagator will become free from all the non-physical parameters of the theory in any of its possible solutions (not only in the iteration one).
Eq. (5.34) shows that gluon may acquire mass dynamically if the denominator of this equation becomes zero at some point $q^{2}=M^{2}$. The expression for the free massive vector particle propagator follows in a rather simple way, indeed. For this goal it is necessary to omit the corresponding invariant function and identify the tadpole term with the gluon mass $M^{2}$. In Minkowski metric (by replacing $q^{2}\rightarrow-q^{2}$ and $\delta_{\mu\nu}\rightarrow g_{\mu\nu}$), one gets
$D^{0}_{\mu\nu}(q)={-i\over(q^{2}-M^{2})}\left[g_{\mu\nu}-{(1-\lambda^{-1})q_{\mu}q_{\nu}\over(q^{2}-\lambda^{-1}M^{2})}\right],$ (7.46)
i.e., it will be expressed in the Stueckelberg gauge [4]. Contracting this expression with respect to $q_{\mu}$ and $q_{\nu}$, one obtains the ST identity (5.35) by replacing there $q^{2}\rightarrow-q^{2}$. In other words, the Stueckelberg gauge (7.46) is a particular case of the generalized gauge-fixing function (5.33). The PT QCD full gluon propagator (6.38) cannot generate the gluon mass at some point, since it remains always massless. The full gluon propagator (5.34) provides a rather convenient framework in order to formulate and further develop the NP renormalization program for the massive gluon fields. This program will suppose to show that the full gluon propagator (5.34), and hence its free massive counterpart (7.46), can not be put on the mass-shell $q^{2}=M^{2}_{g}$, where $M^{2}_{g}$ is the finite gluon pole mass. Acquiring mass dynamically, the gluon, nevertheless, will remain the off-mass-shell object. So that, it will not appear in the physical spectrum. Some aspects of the massive gluon field’s renormalization are discussed in appendix C.
In the presence of the mass squared scale parameter - the tadpole term - the role of the QCD coupling constant $g^{2}$ becomes unimportant. This is also evidence of the ’dimensional transmutation’, $g^{2}\rightarrow\Delta^{2}_{t}(D)$ [2, 36, 37], which occurs whenever a massless theory acquires mass dynamically. It is a general feature of spontaneously symmetry breaking in field theories. We distinguish between the PT/particular/conventional and general solutions to QCD by the explicit presence of the tadpole term in the latter one, and not by the magnitude of the coupling constant even at the regularized gluon fields level yet. In the both cases the gluon fields remain strongly interacted, apart from the AF regime [38, 39]. So that, the general solution to QCD, investigated in section 5 is possible to call as the NP QCD as well.
Any deviation of the full gluon propagator from the free one in QCD requires the presence of a mass squared scale parameter. Even in the AF regime $q^{2}\rightarrow\infty$ there is a scale breaking,
$D_{\mu\nu}(q)\sim\delta_{\mu\nu}\Bigl{[}{ig^{2}\over 1+g^{2}b_{0}\ln(q^{2}/\Lambda^{2}_{QCD})}\Bigr{]}(1/q^{2}),$ (7.47)
@@
which has been discovered many years ago [37, 38]. When $q^{2}\rightarrow\infty$ the dependence on the coupling constant $g^{2}$ vanishes, indeed. The well-known expression (7.47) is written down in the t’ Hooft-Feynman gauge and in the approximation of the main PT logarithms with the colour group factor $b_{0}>0$. $\Lambda^{2}_{QCD}$ determines the scale of the non-trivial PT dynamics in the QCD. However, in section 6, we have already shown that the PT cannot generate the scale parameter, having the dimension of a mass squared. From the renormalization group equations analysis, it follows that any mass to which can be assigned any physical meaning is exponentially suppressed in the weak coupling limit to every order of the PT [2, 36], namely
@@
where $\mu$ is the arbitrary renormalization point. None a finite mass can survive in the PT weak coupling limit or, equivalently, in the PT $q^{2}\rightarrow\infty$ regime, indeed. So the question where the finite mass comes from? cannot be answered by the PT! It is entirely of the NP origin! It has to come from the IR region which is controlled by the tadpole term, as discussed just above. Its finite version should be somehow related to $\Lambda^{2}_{QCD}$ (reminding that the ratio $(\Delta^{2}_{t}(D)/q^{2})$ may appear under the PT logarithms). The explanation of the AF phenomenon (i.e., a scale breaking in the PT regime) will be given when the NP renormalization program of the singular gluon field configurations will be performed.
8 Conclusions
We have derived the novel exact constraint (3.15) on any solution to QCD. For this aim the ST identities for the gauge particles Green’s functions (propagators), complemented by the satisfied transverse relations for the quark and gluon contributions into the full gluon self-energy, have been used. The first of its solutions is called the general solution to QCD and given by the expression (3.16), which formulates the generalized gauge-fixing function approach to QCD. The existence of a new general solution to QCD makes it possible to prove in the most general and unique mathematical ways that the $SU(3)$ colour gauge symmetry of the QCD Lagrangian is broken in its ground state. However, the PT renormalizability will not be affected within the framework of the full gluon propagator. In the PT $q^{2}\rightarrow\infty$ regime it will have the PT renormalizable behaviour $\sim 1/q^{2}$.
The dynamical source of the Lagrangian’s gauge symmetry breakdown in the ground state has to be identified with the constant tadpole term, having the dimension of a mass squared. It is dynamically generated by the self-interaction of the multiple massless gluon modes involving only the point-like 4-gluon vertex, see eq. (2.4) and Fig. 1. The full 4-gluon vertex survives when all the gluon momenta involved go to zero, becoming thus its point-like counterpart, while the full 3-gluon vertex does not survive in this limit. Just this shows the importance of the point-like 4-gluon vertex in the deep IR structure [10, 23, 40] of the theory. By the design the tadpole term is present in the QCD ground state from the very beginning, i.e., not introducing by hand. It is only one which explicitly contributes into the skeleton loops decomposition of the full gluon propagator (shown in Fig. 1), while all the other tadpole-like terms don’t. Unlike them, being the QD but already regularized constant, the tadpole term is not connected to any external gluon momentum, and thus may generate a mass squared scale parameter [6], indeed. Thus, there is no need to extract such a scale parameter in the QCD vacuum in a some other way. In [10] the importance of the tadpole term has been recognized and exploited on an intuitive basis rather than a logical one. However, from now on its existence cannot be undermined. Besides its combinatorial (mathematical) meaning it has a deep dynamical one as a source of the gauge symmetry violation in the QCD vacuum. Its renormalized (i.e., finite) version will determine a scale of the NP dynamics in QCD and in its true ground state.
Due to these impacts our advance findings are:
a. The novel general constraint (3.15) on any solution to QCD has been exactly derived before any proper subtraction scheme has been introduced. The importance of the ST identities and the corresponding satisfied transverse relations has been explicitly shown.
b. Its exact solution (3.16) determines the generalized gauge-fixing function approach to QCD.
c. The proper subtraction scheme has been formulated to remove from the theory the QD constants, apart from the tadpole term, on the general mathematical basis.
d. The system of the relations (5.36) are exact mathematical results, i.e., they are not prescriptions.
e. The new NP approach to the YM theory and its ground state has been formulated, based on the regularized system of eqs.  (5.34)-(5.35). It is defined up to the invariant function with the known asymptotic properties.
f. This system has a correct the PT limit, and is suitable for the NP renormalization program to be performed. It is called as the general solution to QCD, and just it resolves the dilemma formulated at the end of section 2.
In obtaining all these results no any truncations, approximations and assumptions have been made in dealing with the regularized skeleton loop integrals, contributing to the gluon self-energy. The exact analytic derivations have been done, based on a tensor algebra rigorous rules.
The second of the above-mentioned solutions is called the particular solution to QCD, and determined by the relations $\Delta^{2}_{t}(D)=0$ and $\xi=\xi_{0}$ in the exact constraint (3.15). This solution is nothing else but the well-known PT or, equivalently, conventional QCD, which has been studied last five decades. We confirm that the second solution preserves the gauge symmetry of the QCD Lagrangian in its ground state. Both equalities, shown above, are simply prescriptions, i.e., they have to be put valid by hand, as it has been discussed in section 6. In this connection, let us remind that for the tadpole term does not exist the independent mathematical relation according to which it has to be omitted from the theory, while such kind of the relations exist for the the quark and gluon QD regularized constants (see section 4). That is why the tadpole term has to be put zero by hand for the particular solution, which comes from the solution of the exact constraint (3.15), and see also discussion in appendix B.
We can registered the existence of the two different types of the gluon field configurations in the QCD vacuum: the general one and its particular counterpart with the gauge symmetry broken and unbroken, respectively.
The just derived general solution to QCD with a gauge symmetry breaking, but maintaining its PT renormalizability, will apparently satisfy the boundary conditions in QCD at the fundamental quark-gluon level: the confinement at large distances ($q^{2}\rightarrow 0$) and the AF at short distances ($q^{2}\rightarrow\infty$), as discussed in section 7. The particular or, equivalently, the PT/conventional QCD solution with a gauge symmetry preservation cannot satisfy the boundary condition at large distances. It shows the agreement with the AF regime at short distances, but cannot explain it, see discussion in section 7 as well.

### v4

@@ 7 Conventional QCD
From above and now on in this section we denote the corresponding full gluon propagator $D$ by $D^{PT}$ (for the explanation see the text below). Obviously, the expression for the PT full gluon propagator (7.42) describes the propagation of the PT massless gluons, since it has the PT singularity on the mass-shell $q^{2}=0$, i.e., the singularity of the free gluon propagator $\sim 1/q^{2}$. The gluon SD equation (7.42) does not provide any hint how to prevent the free and the PT massless gluon states to appear at large distance ($q^{2}\rightarrow 0$), i.e., this theory is not confining, as emphasized above. Let us note in advance that the equality $\xi=\xi_{0}=\lambda^{-1}$ takes place only for the regularized massless gluon fields. For their renormalized counterparts these gauge-fixing parameters differ by the corresponding renormalization constant.
The conventional or, equivalently, the PT QCD is a well-elaborated quantum gauge field theory [2, 3, 4, 5, 6, 7, 8, 9], and for its framework of the corresponding dynamical equations of motion see recent review [34] (and references therein). Just it has been investigated in full details the last five decades since its formulation in [1]. It is necessary to remind that the coupling constant remains strong, so to call it as $D^{PT}$ is only a mere convention. This theory is not our concern. We present it here within our formalism for the readers convenience to directly compare it with the general solution to QCD. For this aim it is useful to reproduce the system of the relations (5.36)-(5.37) for the PT QCD. It will look like
@@
The first equalities in these relations coincide with the corresponding equalities in the relations (5.36)-(5.37), i.e., in both theories the general and the PT QCD they are exact mathematical results. At the same time, the second equalities in these relations remain prescriptions, since they have to be put zero by hand, as emphasized above.
The dimensional regularization method (DRM) [4, 6, 27, 28, 29], formulated within the framework of the conventional QCD, provides a gauge-invariant scheme to correctly calculate the finite parts of the generally QD loop integrals, while omitting their QD constants, i.e., simply to ignore them and to deal further only with the logarithmic divergences of the PT. This was a prescription rather than an exact result, as pointed out in [4]. However, from now on this prescription has been put on a firm mathematical ground. How to understand such kind of equalities has been already explained in section 5.
8 Confinement of the free gluon states
For the sake of transparency and simplicity, we use the free gluon propagator in the t’ Hooft-Feynman gauge $\lambda^{-1}=1$, i.e., $D^{0}_{\mu\nu}(q)=i{\delta_{\mu\nu}/q^{2}}$. Then the full gluon propagator (5.28), on account of the definitions (5.29) and (5.30), up to the first iteration term looks like
@@ 8 Confinement of the free gluon states
where the terms of order $O_{\mu\nu}(1/q^{2})$ are not shown, for simplicity. In this iteration expansion is explicitly shown only singular part of the corresponding Laurent series, and thus it represents the summation of all the severe IR singularities (more singular than the PT IR singularity $\sim 1/q^{2}$), which can be accumulated by the full gluon propagator. Evidently, such a summation is only possible due to the existence of the tadpole term. The dependence of the coefficients functions $\Phi_{k}(g^{2},...)$ on the coupling constant is not known, i.e., such kind of series are the cluster expansions, mentioned above [23]. These coefficients, being the sum of the infinite number of terms by themselves, are resulted from the skeleton loop integrals, contributing into the full gluon self-energy through the invariant function $\Pi(q^{2};D)$. Let us note here that eq. (8.49) can be expressed in terms of the mass gap through the relation $\Delta^{2}_{t}(D)=Z_{\Delta}\Delta^{2}$ and further including $Z_{\Delta}$ into the above-mentioned coefficient functions. Evidently, the iteration ab  infinity leads to the essential singularity at $q^{2}\rightarrow 0$! in the full gluon propagator (8.49), which will remain massless, but depending on the tadpole term (scale breaking in the IR region). Let us note that the same series can be obtained by expanding the initial NL transcendental relation (5.31) in powers of the ratio $(\Delta^{2}_{t}(D)/q^{2})$.
@@
From the NL iteration expressions (8.48)-(8.49) it follows that the full gluon propagator in the IR region ($q^{2}\rightarrow 0$) is dominated by the tadpole term contribution, while in the UV region ($q^{2}\rightarrow\infty$) it will be dominated by the free gluon one. This is completely different to QED, where the dressed photon propagator behaves at small photon momentum as its free counterpart ($\sim 1/q^{2}$). This is also true for the full gluon propagator in the conventional QCD, investigated in section 6. The principal difference between these theories is due to the existence of the tadpole term in the QCD ground state, which precisely controls the structure of a single full gluon propagator at small $q^{2}$. That is why the dressed PT gluon (7.42) and free gluon states can not appear in the physical spectrum at large distances ($q^{2}\rightarrow 0$), i.e., they will be suppressed in this region by the tadpole term. If this statement will survive the corresponding renormalization program beyond the PT then confinement of the PT and free gluon states will be explained. This is important since there is no any other dynamical mechanism to confine these gluon states to the vacuum of QCD or to keep them inside hadrons. To prove the confinement of the singular full gluon propagator itself due to the above-mentioned essential singularity, though possible, but requires completely different investigation within the theory of distributions [26]. It is the subject of the future work. Let us note that by the NP program we mean that the full gluon propagator will become free from all the non-physical parameters (like the coefficient functions mentioned above) of the theory in any of its possible solutions (not only in the iteration one).
9 Asymptotic freedom
let us now investigate the PT $q^{2}\rightarrow\infty$ limit of the full gluon propagator (5.34). For this aim it is enough to neglect the ratio $(\Delta^{2}_{t}(D)/q^{2})$ in the invariant function eq. (5.31). Also, to replace the the invariant function $\Pi(q^{2})$ by its to leading order logarithm, and to introduce the following relation $d(q^{2})=\alpha_{s}(q^{2};\Lambda^{2})/\alpha_{s}(\lambda)$, where $\Lambda^{2}$ and $\lambda$ are the UV regulating parameter and its dimensionless counterpart, respectively. Then eq. (5.34) will look like
$D_{\mu\nu}(q)\sim iT_{\mu\nu}(q)\Bigl{[}{\alpha_{s}(\lambda)\over 1+\alpha_{s}(\lambda)b_{0}\ln(q^{2}/\Lambda^{2})}\Bigr{]}{1\over q^{2}}+...,$ (9.50)
where $b_{0}>0$ is the colour group factor and $\alpha_{s}(\lambda)$ is the coupling constant squared. Here and below the trivial longitudinal part is not shown, for simplicity. Since we already know that the finite mass gap exists, then without loosing generality, one can put $\Lambda^{2}=f(\lambda)\Delta^{2}=f(\lambda)a^{-1}a\Delta^{2}=f^{\prime}(\lambda)\Lambda^{2}_{QCD}$ with $\Lambda^{2}_{QCD}=a\Delta^{2}$, where $a$ is a some finite number because one cannot identify directly the mass gap $\Delta^{2}$ with $\Lambda^{2}_{QCD}$. The dimensionless functions $f^{\prime}(\lambda)=a^{-1}f(\lambda)$ behave like $\sim\lambda^{2}$ in the $\lambda\rightarrow\infty$ limit, otherwise remaining arbitrary. Doing some algebra, one obtains
$D_{\mu\nu}(q)\sim iT_{\mu\nu}(q)\Bigl{[}{\alpha_{s}\over 1+\alpha_{s}b_{0}\ln(q^{2}/\Lambda^{2}_{QCD})}\Bigr{]}{1\over q^{2}}+...,$ (9.51)
where finite $\alpha_{s}=\alpha_{s}(M_{Z})=0.1184$ is determined by the relation connecting $\alpha_{s}(\lambda)$ to $\ln f^{\prime}(\lambda)$. Eq. (9.51) has been discovered many years ago [38, 39]. In the PT $q^{2}\rightarrow\infty$ limit the dependence on the coupling constant $\alpha_{s}$ vanishes, indeed. The well-known expression (9.51) is written down in the so-called approximation of the main PT logarithms and $\Lambda^{2}_{QCD}$ determines the scale of the non-trivial PT dynamics in the QCD.
@@ 9 Asymptotic freedom
However, in section 6, we have already shown that the PT QCD cannot generate the scale parameter, having the dimension of a mass squared. From the renormalization group equations analysis, it follows that any mass to which can be assigned any physical meaning is exponentially suppressed in the weak coupling limit to every order of the PT [2, 36], namely
@@
where $\mu$ is the arbitrary renormalization point. None a finite mass can survive in the PT weak coupling limit or, equivalently, in the PT $q^{2}\rightarrow\infty$ regime, indeed. So the question where the finite mass comes from? cannot be answered by the PT! It is entirely of the NP origin! It has to come from the IR region which is controlled by the tadpole term, as discussed above. So that, the scale breaking appears under the PT logarithm, while the gluon itself remains massless like in eq. (8.49), where the scale breaking appears in the NL way. In the both cases the scale breaking masses cannot be directly identified with the mass gap, defined in section 6. It should be multiplied by the finite numbers different in each case, in order to get scale breaking in the IR and UV regions. Its exact derivation, and thus the explanation of the AF phenomenon (i.e., a scale breaking in the PT regime) will be given when the corresponding NP renormalization program for the full gluon propagator (5.34) will be performed.
Concluding, all this can be manifestation that ”the problems encountered in perturbation theory are not mere mathematical artefacts but rather signify deep properties of the full theory” [43]. The message we are trying to convey is that the AF phenomenon is the first direct evidence of the validity of the general QCD. In our opinion, the gluon confinement and AF are the two boundary conditions at low- and high-energies in the QCD and its vacuum, to which the general QCD satisfies, while the conventional QCD does not. For these features of the general QCD, one can call it as the confining QCD theory as well. Finally, it is worth emphasizing once more that in the both asymptotics of the iteration/singular solution to the confining QCD, the gluons always remain massless. So that, the QCD Lagrangian’s gauge symmetry is not violated by it, but the mass gap is needed in order to provide its scaling properties, as described above.
10 Discussion and outlook
The general QCD requires the explicit presence of the mass squared scale parameter - the tadpole term - in the QCD ground state, i.e., at the fundamental quark-gluon level, as this was exactly proven in section 4 and used in section 5. In its presence the role of the QCD coupling constant $g^{2}$ becomes unimportant. This is also evidence of the ’dimensional transmutation’, $g^{2}\rightarrow\Delta^{2}_{t}(D)$ [2, 36, 37], which occurs whenever a massless theory acquires mass dynamically. It is a general feature of spontaneously symmetry breaking in field theories. We distinguish between the PT/particular/conventional and general QCD by the explicit presence of the tadpole term in the latter one, and not by the magnitude of the coupling constant even at the regularized gluon fields level yet. In the both cases the gluon fields remain strongly interacted, apart from the AF regime [38, 39].
The general character of our results is convenient to summarize in connection with the Jaffe-Witten’s (JW) [23] theorem, formulated as follows:
JW’s theorem: Yang-Mills existence and Mass Gap. Prove that for any compact simple gauge group $G$, a non-trivial quantum Yang-Mills theory exists on $R^{4}$ and has a mass gap $\Delta>0$. Existence includes establishing some strong axiomatic properties for the Euclidean Green’s functions.
In their description of the theorem it is explained why a correct QCD theory must have the following three properties: 1). The nuclear force is strong but short-ranged due to the existence of the ”mass gap”, a scale parameter $\Delta>0$. 2). The physical particle states are $SU(3)$ invariant,i.e., confinement of colour degrees of freedom at the fundamental quark-gluon level. 3). The chiral symmetry breaking to account for the ”current algebra” theory of soft pions at the hadronic level.
The definition of the mass gap given in [23] as the solution of the Hamiltonian problem that every excitation of the vacuum has energy at least $\Delta>0$ is the mass gap responsible for the second PCAC phase transition at the hadronic level (in our opinion). For example, the glueball state with the most lowest mass, though none of such state has been discovered yet (but this is not a problem, remanding that the Higgs particle [41] has been experimentally found after many years of its prediction). The literary meaning/definition of the mass gap, mentioned above, may appear at the fundamental quark-gluon level as well in the general QCD only in its massive solution for the full gluon propagator (5.34). The gluon may acquire mass dynamically, indeed, if the denominator of this equation becomes zero at some finite point $q^{2}=M^{2}=\Delta^{2}_{t}(D)$. This solution is completely different case from the singular solution, and will be subject for the independent consideration.
The request for a mass gap in the general QCD, can be also summarized like the JW theorem:
Theorem I: Mass Gap existence and Yang-Mills theory. If a non-trivial quantum Yang-Mills theory with gauge group $SU(3)$ exists on $R^{4}$ then it has a mass gap $\Delta^{2}>0$, responsible for the first confinement phase transition in QCD. Existence includes establishing the confining QCD theory, which remaining the PT renormalizable.
In connection with our theorem, let us note that the tadpole term cannot be removed from the QCD theory and its vacuum by any means, apart from putting it zero by hand in order to get to the conventional QCD. Its renormalized (finite) version survives the NP renormalization and thus can be assigned a physical meaning as a scale parameter responsible for the NP effects in the QCD, leading to the formulation of the confining QCD theory. Just the existence of a such scale, conventionally called as a mass gap, provides the suppression of the PT and free gluon states to appear at large distances, i.e., in the physical spectrum, as explained above. It is dynamically generated by the self-interaction of the multiplied massless gluon modes, dominated by the point-like four-gluon vertex [10, 23, 40], i.e., it is a fundamental quantity (”mass without mass” [44]), so that it cannot be a bound state of anything. In [10] the importance of the mass gap has been recognized and exploited on an intuitive basis rather than a logical one, but from now on its existence cannot be undermined.
A few additional clarifying remarks are in order. Let us remind that the present QCD theory, i.e., conventional one in our terminology, possesses none of the above-mentioned three necessary properties required by JW for QCD to be successful theory of the strong interactions at the both microscopic and macroscopic levels. Unlike this theory, the confining QCD theory, discovered in this work, possesses at least the two of those properties. It has a characteristic mass scale parameter even at the fundamental quark-gluon level. Precisely, the existence of this scale in the theory confines the colour gluons to the QCD ground state. We strongly believe that quark confinement and ”chiral symmetry breaking” will be also confirmed after solution of the quark full SD equation by the substitution of the full gluon propagator, determined within the confining QCD theory. It is worth emphasize once more that we have explicitly shown that the first confinement phase transition in QCD should have its own finite scale parameter $\Delta^{2}$, reflected in our Theorem I. In order to describe the scales of the different NP effects, such as the scales breaking, the existence of the string tension, etc., it should undergo the finite transformations, i.e., to be multiplied by the different finite constants.
Taking into account that the JW theorem has been formulated in the most general terms, absorbing the different interpretations/meanings/understanding of a mass gap itself, it was quite natural and suitable for us to adopt the JW mass gap name for our characteristic mass scale parameter. In other words, we understand the mass gap conception in more wider context than its definition given by JW. Nevertheless, to summarize all our advance results in terms of the theorem similar to the JW theorem, makes sense and is very useful for further development of the correct (confining) QCD theory and understanding of its true vacuum dynamical and gauge structures. In obtaining them no any truncations, approximations and assumptions have been made in dealing with the regularized skeleton loop integrals, contributing into the full gluon self-energy. The exact analytical derivations have been done, based on a tensor algebra rigorous rules only.